The equation of the normal to the curve y=x^3+e^x at x=0 is

Grade:11

1 Answers

ROSHAN MUJEEB
3 years ago
Dear Student
Different the equation wrt y. -dx/dy is the slope of normal.
Differentiating wrt y gives 1=3x^2dx/dy + e^xdx/dy. Put x=0
It gives 1=dx/dy . Slope of normal is -1. The coordinate is (0,1)
equation of line is y-1= -x. Hope you got it